OpenStudy (anonymous):

A tortoise makes a journey in two parts: it can either walk at 4cm/s or crawl at 3cm/s. If the tortoise walks the first part and crawls the second, it takes 110 seconds. If it crawls the first part and walks the second, it takes 100 seconds.Find the lengths of the two parts of the journey.

OpenStudy (bahrom7893):

time = distance/speed Let a be the first half and b be the second half. Then: a/4 + b/3 = 110 a/3 + b/4 = 100 Find a and b

OpenStudy (bahrom7893):

Can you finish the rest?

OpenStudy (bahrom7893):

HELLLOOOO? Are you there?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i m trying to solve it

OpenStudy (bahrom7893):

ok

OpenStudy (anonymous):

but i got stuck in fractions

OpenStudy (bahrom7893):

a/4 + b/3=110 a + 4b/3 = 440 a = 440 - 4b/3

OpenStudy (bahrom7893):

(440 - 4b/3 )/3+b/4 = 100 4(440-4b/3) + 3b = 1200 1760 - 16b/3 + 3b = 1200 5280 - 16b + 9b = 3600 -7b = -1680 b = 240

OpenStudy (bahrom7893):

a = 440 - 4b/3 = 440 - 4*240/3 = 440 - 4*80 = 440 - 320 = 120

OpenStudy (anonymous):

thnx for helping me